OpenStudy (anonymous):

I need help with this 2^(Log_2_(3) - Log_2_(5))

OpenStudy (ash2326):

Is this your question \[2^{\log_{2} 3-\log_{2}5}\]?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

we know \[ \log_{a} b -\log_{a} c= \log_{a} {\frac{b}{c}}\] so her we'll get \[ 2^{log_{2} 3-log_{2} 5}= 2^{log_{2} \frac{3}{5}}\] now we know that \[a^{\log_{a} b}= b\] so we get \[2^{log_{2} \frac{3}{5}}= \frac{3}{5}\]

OpenStudy (anonymous):

ah ah!